There is a coffee shop selling coffee and bagels. There are three types of consumers. Type A’s WTP (willingness to pay) for coffee is 5, and WTP for bagel is 0. Type B’s WTP for coffee is 0, and WTP for bagel is 5. Type C’s WTP for coffee is 4, and WTP for bagel is 4. Assume there are as many A types as B and C types. The coffee shop is trying to maximize revenue. Which statement is correct?

(a) It should sell the coffee and bagel only as a bundle together at a price of 5.

(b) It should sell the coffee and bagel separately each at a price of 4.

(c) It should sell the coffee and bagel separately each at a price of 4 and also as a bundle at 5.

(d) It should sell the coffee and bagel separately each at a price of 5 and also as a bundle at 8.

The answer is d not a, b, or c
Answer #1

If the coffee shop choses to sell coffee and bagel as a bundle at price of 5 then Type A will purchase but as far as Type C is considered, he will not purchase the product because his willingness to pay total is 8 (4 = coffee, 4= bagel). Marginal revenue=4.

If the coffee shop choses to sell the coffee and bagel seperately each at a price of 4 then Type C will only purchase , rest Type A and Type B will not purchase because their WTP for coffee is 5 for typeA and 0 for Type B whereas coffee shop is selling at 4 which is lower than what actually they are willing to pay. Same goes for bagel typeA willing to pay 0 but typeB willing to pay 5 but firm is willing to Sell at 4. This does not satisfy in order to increase MR. Marginal revenue=8

Is the coffee shop choses to sell seperately coffee and bagel at price of 4 and bundle of 5. All types of consumers will be satisfied but marginal revenue will not be more as compared to option D. Seperately at price 4, 4+4=8 and bundle at price 5 which is lower than option D . So total marginal revenue= 4+4+5=13

If coffee shop choses to sel seeprately coffee and bagel at price 5 as bundle at price 8. Then here marginal revenue will be more because seperately at proce 5 will give revenue 5+5=10 and Marginal revenue will be higher from all options=8. Total marginal revenue= 5+5+8=18

So highest marginal revenue is for point d.